Penelope spends half her income on housing... And budgets the other half for the following: 1/4 on clothes, 1/3 on entertainment, and 1/4 on transportation. She saves the remaining $40. What is her monthly income?
Answer by lwsshak3(11628)

Penelope spends half her income on housing... And budgets the other half for the following: 1/4 on clothes, 1/3 on entertainment, and 1/4 on transportation. She saves the remaining $40. What is her monthly income?
**
let x=Penelope's monthly income
x/2+x/2(1/4+1/3+1/4)+40=x
x/2+x/2(10/12)+40=x
x/2+5x/12+40=x
6x/12+5x/12+40=x
11x/12+40=x=12x/12
12x/12-11x/12=40
x/12=40
x=480
ans:
Penelope's monthly income=$480